OpenStudy (anonymous):

2cos2n@+1/2cos@+1=(2cos@-1)(2cos@-1)(2cos2square@-1).. (2cos2n-1@-1).prove it..

OpenStudy (anonymous):

having a hard time understanding, is this the question: \[\frac{2\cos(2n\theta+1)}{2\cos(\theta)} = \cdots\] im not even sure what the right hand side is....

OpenStudy (anonymous):

oh oh, maybe: \[2\cos^{2n}(\theta+1) \] instead?

myininaya (myininaya):

hey shruti can you write this using the equation thingy below if you want you can use frac{somethin}{something} to enter in fractions

OpenStudy (anonymous):

or maybe scan or take a picture and upload it.

myininaya (myininaya):

that might be easier

OpenStudy (anonymous):

dont know what it is exactly , but pretty good chance it involves complex numbers and sum of a gp